# Drainage and lifetime of thin liquid films: the role of salinity and convective evaporation

### v1

@@ 5.1 Mean lifetime and convective evaporation
To investigate film lifetime, we repeat the drainage experiment at least a hundred times to obtain robust statistics. In the cases with large lifetime fluctuations (relative humidity $\mathcal{R}_{H}>70\%$) the experiment is repeated two hundred times. Given the limited influence of the solution on drainage, we use solutions of Triton X-100 without any salt or glycerol added to focus on the effect of the air around the film through humidity or flow. In particular, we investigate the dependence of lifetime with humidity because of the stark contrast between its lack of effect on drainage and the key role it plays in determining the lifetime of soap films. When the experiment is repeated, a difficulty arises: sometimes the lifetime of the film for a given set of conditions varies drastically throughout repetitions. This difficulty, which was previously noted by Poulain et al., (2018) in the case of surface bubbles is particularly striking at humidities close to saturation (Fig. 5 (a), 90% relative humidity). The distribution of lifetimes appears to change drastically and irreversibly towards extremely long times (tens of minutes) even though the experimental conditions do not vary (the temperature and relative humidity are continuously recorded in the chamber, see Fig. 5 (a)).
Figure 5: (a) Example of a time series of lifetime measurements without the convection activated in the chamber. (top) Relative humidity (blue) and temperature (orange) close to the film. (bottom) Lifetime of the films over the course of the experiment.
@@
These modulations in the distribution of lifetimes are a major source of uncertainty and difficulties in drawing conclusion in studies about surface bubbles. We interpret these fluctuations as being due to the evaporation of the surface from which the film is drawn. As the surface evaporates, the air right above it gets charged in water vapor, locally increasing the humidity. The humidity close to the film is therefore closer to saturation than what would be expected from the far field humidity measurement that controls the humidity in the chamber. Humid air being lighter than dry air, a plume of high humidity air forms over the center of the bath (Dollet and Boulogne,, 2017; Boulogne and Dollet,, 2018). This plume has not been characterized experimentally, however plumes around evaporating droplets are common (Dehaeck et al.,, 2014; Dietrich et al.,, 2016) and the mass transfer problem is in this case completely analogous to the heat transfer problem of a heated disk (Lienhard and Lienhard,, 2024) where plumes of warm air have been widely studied (Torrance et al.,, 1969; Lopez and Marques,, 2013; Kwak et al.,, 2018; Khrapunov and Chumakov,, 2020). When we introduce airflow in the chamber, this localized increase in humidity is washed away by the turbulent mixing in the chamber, making the humidity field homogeneous. Using the Rayleigh-Bénard instability (Fig. 1 (b)), the airflow is weak enough such that it does not disturb the drainage studied in the previous section but strong enough to continuously mix and renew the air in contact with the film. We show a comparison of the drainage curves with and without airflow in Fig. 5 (b) and demonstrate that the drainage is not modified.
@@ 5.2 Evaporative flux as a function of relative humidity
with $\rho_{s}$ and $\rho_{\infty}$ the air density close to the film and in the far field, respectively. The length scale $\mathcal{L}$ is in our case of the order of the size of the bath or the film height (both about 1 cm). The Grashof number for this problem is typically $\mathcal{O}(10^{4}-10^{5})$. Previous studies have focused on the mass transfer due to evaporation of the bath alone (Dollet and Boulogne,, 2017) and the film alone (Boulogne and Dollet,, 2018). In the case of the film alone, it is possible to find theoretically a mass transfer coefficient scaling like $Gr^{1/4}$.
@@
Finally, the dependence of the mass transfer coefficient with relative humidity can be put forward: $k_{e}\propto(1-\mathcal{R}_{H})^{1/4}$ in the case of a film without a bath, giving $t_{f}\propto(1-\mathcal{R}_{H})^{-5/4}$. In the case of our problem, we therefore expect $t_{f}\propto(1-\mathcal{R}_{H})^{-(1+\alpha)}$ and find empirically that $\alpha=0.5$ is a good fit of our data (black dotted line of Fig. 6). As a result, the mass transfer coefficient scales like $k_{e}\propto Gr^{0.5}$ in this natural convection case. This scaling is the result of the non-trivial flow generated by the combined evaporation of the bath and the film at the same time. The long term fluctuations visible at high humidity (Fig 5 (a)) are presumably a qualitative change of the flow around the film that also alters the mass transfer coefficient.
In the case where mixing is activated, finding the mass transfer coefficient is simpler as the convection is forced, and the mass transfer coefficient $k_{e}$ is independent of the relative humidity. The mass transfer coefficient is controlled by the turbulent flow and only depends on the velocity resulting from the Rayleigh-Bénard convection $u_{\textsc{rb}}$, i.e. $k_{e}=f(Re,Sc$), with $Re=\rho_{a}u_{\textsc{rb}}W/\mu_{a}$ the Reynolds number associated with the flow over the film, and $Sc=\nu_{a}/\mathcal{D}$ the Schmidt number comparing kinematic viscosity with the mass diffusivity of water vapor $\mathcal{D}$ (Lienhard and Lienhard,, 2024). As a result, we obtain the scaling
@@ 5.3 Evolution of the black film extent
From high speed observations of the film bursting (Fig. 7 (a)), we can notice that the burst location is always at the top: where the film is the thinnest and where the black film is located (Saulnier et al.,, 2014; Pasquet et al.,, 2024). To link our observations on lifetime and drainage, we therefore investigate the time evolution of this region of the film, which usually is small enough to not alter the drainage but appears to play an important role for the lifetime. In particular, we aim to put forward its dependence on humidity and the mixing of the atmosphere in the chamber. Because the thickness of this region is not accessible with our visible light visualization method, we can only probe its spatial extent within the frame. As the film ages, the black film grows and takes up more and more space. For extremely long-lasting films (with a quiescent atmosphere), it may fill almost half the frame. In Fig. 7 (b), we show examples of the evolution of the vertical extent of the black film $z_{\textsc{bf}}$ for each relative humidity condition and with convection in the atmosphere. We measure the vertical extent of the film in the middle where it is the largest. The black film grows continuously throughout the lifetime of the films at a rate that increases with time. In contrast to the drainage, the extent of the black film depends on relative humidity: it grows much faster in a dry atmosphere, and does so from the very beginning of the lifetime of the films. At long times, we find that the evolution of the black film follows closely a power law with an exponent of 2:
$z_{\textsc{bf}}/H=\mathcal{B}\times(t/t_{0})^{2},$ (10)
@@ Appendix A Langmuir through measurements
Every solution used in this study was tested in a Langmuir Through (KSV NIMA, model KN 1003) to record its surface tension isotherm. We record the dynamic properties of the surface by setting aside a small amount of solution and placing it in the through, letting it rest until the surface tension does not evolve significantly. The surface tension is measured with a platinum Wilhelmy plate. After that, two Teflon barriers compress the surface at a rate of 270 mm/min allowing us to obtain the surface tension $\gamma$ as a function of the through area.

### v2

@@ 5 Film lifetime
In this section, we discuss the film lifetime and the effect of convective evaporation on the lifetime statistics and mean values. We rationalize the observations using scaling arguments as well as measurements of the rupture location in the black film and its evolution.
@@ 5.1 Mean lifetime and convective evaporation
To investigate film lifetime, we repeat the drainage experiment at least a hundred times to obtain robust statistics. In the cases with large lifetime fluctuations (relative humidity $\mathcal{R}_{H}>70\%$) the experiment is repeated two hundred times. Given the limited influence of the solution on drainage, we use solutions of Triton X-100 without any salt or glycerol added to focus on the effect of the air around the film through humidity or flow. In particular, we investigate the dependence of lifetime with humidity because of the stark contrast between its lack of effect on drainage and the key role it plays in determining the lifetime of soap films. When the experiment is repeated, a difficulty arises: sometimes the lifetime of the film for a given set of conditions varies drastically throughout repetitions (see Fig. 5 (a), where the mean lifetime changes by a factor five with no variation of the control parameters). This difficulty, which was previously noted by Poulain et al., (2018) in the case of surface bubbles is particularly striking at humidities close to saturation. The distribution of lifetimes appears to change drastically and irreversibly towards extremely long times (tens of minutes) even though the experimental conditions do not vary (the temperature and relative humidity are continuously recorded in the chamber, see Fig. 5 (a)).
Figure 5: (a) Example of a time series of lifetime measurements without the convection activated in the chamber. After about an hour, the lifetime drastically changes while the parameters (temperature and humidity) remain constant. (top) Relative humidity (blue) and temperature (orange) close to the film. (bottom) Lifetime of the films over the course of the experiment.
@@
These modulations in the distribution of lifetimes are a major source of uncertainty and difficulties in drawing conclusion in studies about surface bubbles. We interpret these fluctuations in our case as being due to slow modulations of the evaporation of the surface from which the film is drawn. As the surface evaporates, the air right above it gets charged in water vapor, locally increasing the humidity. The humidity close to the film is therefore closer to saturation than what would be expected from the far field humidity measurement that controls the humidity in the chamber. Exactly how or why this local humidity increase is modulated remains for now uncertain, one possibility being the formation or destabilization of a plume of humid air: as it is lighter than dry air, a plume of high humidity air forms over the center of the bath (Dollet and Boulogne,, 2017; Boulogne and Dollet,, 2018). This plume has not been characterized experimentally, however plumes around evaporating droplets are common (Dehaeck et al.,, 2014; Dietrich et al.,, 2016) and the mass transfer problem is in this case completely analogous to the heat transfer problem of a heated disk (Lienhard and Lienhard,, 2024) where plumes of warm air have been widely studied (Torrance et al.,, 1969; Lopez and Marques,, 2013; Kwak et al.,, 2018; Khrapunov and Chumakov,, 2020). When we introduce airflow in the chamber, this localized increase in humidity is washed away by the turbulent mixing in the chamber, making the humidity field homogeneous. Using the Rayleigh-Bénard instability (Fig. 1 (b)), the airflow is weak enough such that it does not disturb the drainage studied in the previous section but strong enough to continuously mix and renew the air in contact with the film. We show a comparison of the drainage curves with and without airflow in Fig. 5 (b) and demonstrate that the drainage is not modified.
@@ 5.2 Evaporative flux as a function of relative humidity
with $\rho_{s}$ and $\rho_{\infty}$ the air density close to the film (and therefore saturated with humidity) and in the far field, respectively. The length scale $\mathcal{L}$ is in our case of the order of the size of the bath or the film height (both about 1 cm). The Grashof number for this problem is typically $\mathcal{O}(10^{4}-10^{5})$. Previous studies have focused on the mass transfer due to evaporation of the bath alone (Dollet and Boulogne,, 2017) and the film alone (Boulogne and Dollet,, 2018). In the case of the film alone, it is possible to theoretically find a scaling: $k_{e}\propto Gr^{1/4}$.
@@
Finally, the dependence of the mass transfer coefficient with relative humidity can be put forward: $k_{e}\propto(1-\mathcal{R}_{H})^{1/4}$ in the case of a film without a bath, giving $t_{f}\propto(1-\mathcal{R}_{H})^{-5/4}$.
In the case of our problem, because of the geometry combining the evaporation of the bath and the film, we do not know the relation between $k_{e}$ and $Gr$. We make the hypothesis that a similar scaling to the one we found for a film without a bath exists, but with an exponent $\alpha$ that may be different (i.e. $k_{e}\propto Gr^{\alpha}$). We therefore expect $t_{f}\propto(1-\mathcal{R}_{H})^{-(1+\alpha)}$ and find empirically that $\alpha=0.5$ is a good fit of our data (black dotted line of Fig. 6). This scaling is the result of the non-trivial flow generated by the combined evaporation of the bath and the film at the same time.
In the case where mixing is activated, finding the mass transfer coefficient is simpler as the convection is forced. In that case, we make the hypothesis that the mass transfer coefficient $k_{e}$ is independent of the relative humidity and controlled by the turbulent flow, only depending on the velocity resulting from the Rayleigh-Bénard convection $u_{\textsc{rb}}$ (i.e. the evaporation flux associated with natural convection is assumed to be negligible compared to the one associated with forced convection), i.e. $k_{e}=f(Re,Sc$), with $Re=\rho_{a}u_{\textsc{rb}}W/\mu_{a}$ the Reynolds number associated with the flow over the film, and $Sc=\nu_{a}/\mathcal{D}$ the Schmidt number comparing kinematic viscosity with the mass diffusivity of water vapor $\mathcal{D}$ (Lienhard and Lienhard,, 2024). As a result, we obtain the scaling
@@ 5.3 Evolution of the black film extent
From high speed observations of the film bursting (Fig. 7 (a)), we can notice that the burst location is always at the top: where the film is the thinnest and where the black film is located (Saulnier et al.,, 2014; Pasquet et al.,, 2024). To link our observations on lifetime and drainage, we therefore investigate the time evolution of this region of the film, which usually is small enough to not alter the drainage but appears to play an important role for the lifetime. In particular, we aim to put forward its dependence on humidity and the mixing of the atmosphere in the chamber. Because the thickness of this region is not accessible with our visible light visualization method, we can only probe its spatial extent within the frame. As the film ages, the black film grows and takes up more and more space. For extremely long-lasting films (with a quiescent atmosphere), it may fill almost half the frame. In Fig. 7 (b), we show examples of the evolution of the vertical extent of the black film $z_{\textsc{bf}}$ for each relative humidity condition and with convection in the atmosphere. We measure the vertical extent of the film in the middle where it is the largest. The black film grows continuously throughout the lifetime of the films at a rate that increases with time. In contrast to the drainage, the extent of the black film depends on relative humidity: it grows much faster in a dry atmosphere, and does so from the very beginning of the lifetime of the films. At long times, we find empirically that the evolution of the black film follows closely a power law with an exponent of 2:
$z_{\textsc{bf}}/H=\mathcal{B}\times(t/t_{\nu})^{2},$ (9)
@@ 6 Conclusion
Declaration of Interests The authors report no conflict of interest.
@@ Appendix A Langmuir through measurements
Every solution used in this study was tested in a Langmuir trough (KSV NIMA, model KN 1003) using the Wilhelmy plate method to record its surface tension isotherm. We record the dynamic properties of the surface by setting aside a small amount of solution and placing it in the trough, letting it rest until the surface tension does not evolve significantly. The surface tension is measured with a platinum Wilhelmy plate. After that, two Teflon barriers compress the surface at a rate of 270 mm/min allowing us to obtain the surface tension $\gamma$ as a function of the through area.